OpenStudy (anonymous):

What is the exponential function modeled by the following table? x f(x) 2 9 3 27 4 81 f(x) = x3 f(x) = 3x f(x) = x2 + 5 f(x) = 2x + 5 @$w3G_Godd

OpenStudy (anonymous):

x^3 3^x x^2+5 2^x+5

OpenStudy (anonymous):

Just plug in x value in each choice and see which one gives you the corresponding f(x). Also the x is exponent in all the choices.

OpenStudy (anonymous):

I got B

OpenStudy (anonymous):

In what direction and by how many units is the graph of f(x) = 6 sin(2x + π) − 5 vertically and horizontally shifted? Down 5, left pi over 2 Up 5, left pi over 2 Down 5, right pi over 2 Up 5, right pi over 2

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

k im back

OpenStudy (anonymous):

i actually dont know the answer to this one

OpenStudy (anonymous):

What is the x-coordinate for the minimum point in the function f(x) = 4 cos(2x − π) from x = 0 to x < 2π? x = pi over 2 x = pi over 4 x = 2π x = π

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the logarithmic function modeled by the following table? x f(x) 9 2 27 3 81 4 f(x) = logx3 f(x) = 3 log10x f(x) = x log103 f(x) = log3x

OpenStudy (anonymous):

the answer is f(x)=log3x Do u need explanation ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What function best represents a sine function with an amplitude of 4, a period of pi over 6, and a midline at y = −2? f(x) = −2 sin(x − pi over 6) + 4 f(x) = 4 sin 12x − 2 f(x) = 4 sin(x − pi over 6) − 2 f(x) = −2 sin 12x + 4